OpenStudy (anonymous):

simplify √5(10-4√2) A) 14 B) 15√2 C) 5√2-4√10 D) None of the above

OpenStudy (anonymous):

C

OpenStudy (anonymous):

bo its none of the above

OpenStudy (anonymous):

no*

jimthompson5910 (jim_thompson5910):

Did you mean \[\Large \sqrt{5}(\sqrt{10}-4\sqrt{2})\] ???

OpenStudy (triciaal):

that person is not here since yesterday

OpenStudy (anonymous):

so 10-4 is 6 then it becomes 6 - sqrt2 sqrt 5 (6 - sqrt 2) 6 x sqrt 5 - sqrt 5 x sqrt 2 =( 6 x sqrt 5) (sqrt 5 x sqrt 2) = (6 x sqrt 5) - sqrt 10